# Scaling Properties of Diffusion Models for Perceptual Tasks

### v1

@@ Abstract
In this paper, we argue that iterative computation with diffusion models offers a powerful paradigm for not only generation but also visual perception tasks. We unify tasks such as depth estimation, optical flow, and segmentation under image-to-image translation, and show how diffusion models benefit from scaling training and test-time compute for these perception tasks. Through a careful analysis of these scaling behaviors, we present various techniques to efficiently train diffusion models for visual perception tasks. Our models achieve improved or comparable performance to state-of-the-art methods using significantly less data and compute. We release code and models at scaling-diffusion-perception.github.io.
@@ 1 Introduction
Diffusion and autoregressive models have emerged as powerful techniques for generating images and videos, while showing excellent scaling behaviors. One could attribute part of the success of diffusion models to iterative computation. At inference, diffusion models perform iterative denoising, consuming approximately one to two orders of magnitude more compute compared to a single forward pass.
In this work, we explore how to efficiently scale diffusion model training and inference for downstream perceptual tasks (inverse problems). Marigold (Ke et al., 2024) and FlowDiffuser (Luo et al., 2024) partially answered this question for depth estimation and optical flow individually, showing that diffusion models are indeed suited for such perceptual tasks. Our work explores how to unify a wide range of perceptual tasks, from low-level optical flow to mid-level depth estimation and high-level semantic segmentation / occlusion reasoning, through a detailed study on train/test-time compute scaling behavior.
To study the scaling properties of diffusion models for inverse problems both at training and test-time, we pre-train various dense and mixture-of-expert models for class-conditional image generation. Then, we fine-tune these pre-trained models on downstream perception tasks. We study scaling properties of fine-tuning by varying model size, data resolution, and pre-training compute. In addition, we apply efficient training strategies such as upcycling dense checkpoints to mixture-of-experts models. We also experiment with various test-time compute scaling techniques, including increasing diffusion steps, test-time ensembling, increasing active model experts, and reorganizing compute with noise variance schedules.
We display the effectiveness of leveraging iterative feedback computation through diffusion models for visual perception tasks with three key contributions:
  • We perform an in-depth study of train/test-time compute scaling laws across all layers of the stack, including pre-training, fine-tuning, and diffusion inference. Specifically, we perform our study on the monocular depth estimation task.
  • We show how to transfer the scaling laws derived for depth estimation to boost performance on tasks such as optical flow or amodal segmentation for both training and inference.
  • We apply all of our scaling strategies to efficiently train a generalist mixture-of-experts model on perception tasks, achieving state-of-the-art results across various benchmarks.
Figure 1: A Unified Framework: We fine-tune a pre-trained Diffusion Model (DM), for visual perception tasks. We take a RGB image, and a conditional image (i.e. next video frame, occlusion mask, etc.), along with the noised image of the ground truth prediction. Our model generates predictions for various visual tasks such as monocular depth estimation, optical flow prediction, and amodal segmentation / completion, based on the conditional task embedding. We train one generalist model that can perform all three tasks with exceptional performance.
@@ 2 Related Work
Generative Modeling: Generative modeling has been studied under various methods, including VAEs (Kingma, 2013), GANs (Goodfellow et al., 2014), Normalizing Flows (Rezende & Mohamed, 2015), Autoregressive models (van den Oord et al., 2016), and Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020). Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) have shown impressive scaling behaviors, becoming the de facto generative tool for many image and video generation models. Notable examples include Latent Diffusion Models (Rombach et al., 2022), which enhanced efficiency by operating in a compressed latent space, Imagen (Saharia et al., 2022), which generates samples in pixel space with increasing resolution, and Consistency Models (Song et al., 2023), which aim to accelerate sampling while maintaining generation quality. Recent methods like Rectified Flow (Liu et al., 2022) and Flow Matching (Lipman et al., 2023) employ training objectives inspired by optimal transport to model continuous vector fields that map data to target distributions, eliminating the discrete steps typical in diffusion models. Rectified Flow mitigates numerical issues in training by applying flow regularization, and Flow Matching offers efficient sampling with fewer discretization artifacts, making them promising alternatives to diffusion-based approaches for high-quality generation. Apart from diffusion models, Parti (Yu et al., 2022) and MARS (He et al., 2024) showcased the potential of autoregressive models for image generation, and the Muse architecture (Chang et al., 2023) introduced a masked image generation approach using transformers.
@@ 6.1 Depth Estimation
We combine our findings from the ablation studies on depth estimation to create a model with the best training and inference configurations. We train a DiT-XL model from (Peebles & Xie, 2023) on depth estimation data from Hypersim for 30K steps with a batch size of 1024, resolution of $512\times 512$, and a learning rate exponentially decaying from $1.2e$-$4$ to $1.2e$-$6$. We use median compilation ensembling with a cosine noise variance schedule. From our scaling experiments, we found the optimal configuration for inference to be 200 denoising steps with $N=5$ samples for ensembling. As shown in Table 3, our model achieves the same validation performance as Marigold on the Hypersim dataset, while being trained with lower resolution images and approximately three orders of magnitude less pre-training data and compute.

### v2

@@ Abstract
In this paper, we argue that iterative computation with diffusion models offers a powerful paradigm for not only generation but also visual perception tasks. We unify tasks such as depth estimation, optical flow, and amodal segmentation under the framework of image-to-image translation, and show how diffusion models benefit from scaling training and test-time compute for these perceptual tasks. Through a careful analysis of these scaling properties, we formulate compute-optimal training and inference recipes to scale diffusion models for visual perception tasks. Our models achieve competitive performance to state-of-the-art methods using significantly less data and compute. We release code and models at scaling-diffusion-perception.github.io.
@@ 1 Introduction
Diffusion models have emerged as powerful techniques for generating images and videos, while showing excellent scaling behaviors. However, diffusion models can also be utilized for downstream prediction for inverse vision problems. In this paper, we present a unified framework to perform depth estimation, optical flow estimation, and amodal segmentation with a single diffusion model, as illustrated in Figure 1.
Previous works such as Marigold (Ke et al., 2024), FlowDiffuser (Luo et al., 2024), and pix2gestalt (Ozguroglu et al., 2024) demonstrate the potential of repurposing image diffusion models for various inverse vision tasks individually. Building on this foundation, we perform an extensive empirical study, establishing scaling power laws for depth estimation, and display their transferability to other perceptual tasks. Using insights from these scaling laws, we formulate compute-optimal recipes for diffusion training and inference. We are the first to show that efficiently scaling compute for diffusion models leads to significant performance gains in downstream perceptual tasks.
Recent works have also focused on scaling test-time compute to enhance the capabilities of modern LLMs, as demonstrated by OpenAI’s o1 model (OpenAI, 2024). As Noam Brown noted, “It turned out that having a bot think for just 20 seconds in a hand of poker got the same boosting performance as scaling up the model by 100,000x and training it for 100,000 times longer.” In our experiments, we realize a similar trade off between allocating more compute during training versus test-time for diffusion models.
We scale test-time compute by exploiting the iterative and stochastic nature of diffusion to increase the number of denoising steps. By allocating more compute to early denoising steps, and ensembling multiple denoised predictions, we consistently achieve higher accuracy on these perceptual tasks. Our results provide evidence of the benefits of scaling test-time compute for inverse vision problems under constrained compute budgets, bringing a new perspective to the conventional paradigm of training-centric scaling for generative models.
Figure 1: A Unified Framework: We fine-tune a pre-trained Diffusion Model (DM), for visual perception tasks. We take a RGB image, and a conditional image (i.e. next video frame, occlusion mask, etc.), along with the noised image of the ground truth prediction. Our model generates predictions for visual tasks such as depth estimation, optical flow prediction, and amodal segmentation, based on the conditional task embedding. We train a generalist model that can perform all three tasks with exceptional performance.
@@ 2 Related Work
Generative Modeling: Generative modeling has been studied under various methods, including VAEs (Kingma, 2013), GANs (Goodfellow et al., 2014), Normalizing Flows (Rezende & Mohamed, 2015), Autoregressive models (van den Oord et al., 2016), and Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020). Denoising Diffusion Probabilistic Models (DDPMs) (Ho et al., 2020) have shown impressive scaling behaviors for many image and video generation models. Notable examples include Latent Diffusion Models (Rombach et al., 2022), which enhanced efficiency by operating in a compressed latent space, Imagen (Saharia et al., 2022), which generates samples in pixel space with increasing resolution, and Consistency Models (Song et al., 2023), which aim to accelerate sampling while maintaining generation quality. Recent methods like Rectified Flow (Liu et al., 2022) and Flow Matching (Lipman et al., 2023) employ training objectives inspired by optimal transport to model continuous vector fields that map data to target distributions, eliminating the discrete formulation of diffusion models. Rectified Flow mitigates numerical issues in training by applying flow regularization, and Flow Matching offers efficient sampling with fewer discretization artifacts, making them promising alternatives to diffusion for high-quality generation. Apart from diffusion models, Parti (Yu et al., 2022) and MARS (He et al., 2024) showcased the potential of autoregressive models for image generation, and the Muse architecture (Chang et al., 2023) introduced a masked image generation approach using transformers.
@@ 6.1 Depth Estimation
We combine our findings from the ablation studies on depth estimation to create a model with the best training and inference configurations. We train a DiT-XL model from (Peebles & Xie, 2023) on depth estimation data from Hypersim for 30K steps with a batch size of 1024, resolution of $512\times 512$, and a learning rate exponentially decaying from $1.2e$-$4$ to $1.2e$-$6$. We use median compilation ensembling with a cosine noise variance schedule. From our scaling experiments, we found the optimal configuration for inference to be 200 denoising steps with $N=5$ samples for ensembling. As shown in Table 3, our model achieves the same validation performance as Marigold on the Hypersim dataset and better performance on the ETH3D test set while being trained with lower resolution images and approximately three orders of magnitude less pre-training data and compute.